# Nocturnal Eye–Inspired Liquid–to-Gas Phase Change Soft Actuator with Laser-Induced-Graphene : Enhanced Environmental Light Harvesting and Photothermal Conversion

### v1

@@ 4.1 Experimental analysis of actuation driven by light On-Off cycling
The experimental setup is illustrated in Fig. 5. The fabricated actuator was positioned 50 mm away from the light source during the experiments. The light bulb specifications include Central luminance is 300  cd, total luminous flux is 630 lm, and beam angle is 60°. An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) was used to capture the actuator deformation, and ImageJ (citation) was employed for analyzing the bending angles and deformation measurements.
Figure 6A and C shows the bending behavior during light irradiation and the subsequent recovery upon light off, along with the corresponding analysis results. When exposed to light, the without-LIG actuator exhibited delayed activation, with a 63% response time of 141.7 seconds and approximately 220 seconds required to reach maximum bending angle. In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds and reached maximum bending angle in approximately 90 seconds, showing an 54.1% improvement in responsiveness compared to the without-LIG actuator.
Regarding the recovery phase after light off (Fig 6B and D, the without-LIG actuator showed a 63% response time of 42.7 seconds and required 120 seconds to return to its initial state. The proposed actuator with LIG demonstrated enhanced recovery characteristics with a 63% response time of 22.1 seconds and approximately 60 seconds to return to its initial state, representing an 48.2% improvement in recovery performance compared to the without-LIG actuator.
Figure 7: Analysis of LIG-transferred soft actuators under different internal and environmental conditions: A-C: When the distance from the light source was constant (50 mm), the deformation ratio (A), plateau angle (B), and time to reach plateau state (C) were measured for internal liquid volumes of 20, 30, and 40 µL. D-F: When the internal liquid volume was constant (30 µL), the deformation ratio (D), plateau angle (E), and time to reach plateau state (F) were measured for distances of 50, 75, and 100 mm between the light source and actuator. The plateau state was defined as the condition where the angular variation did not increase by more than 1° over a 30-second period. The verification was conducted using three actuators (n=3) that exhibited similar plateau angles when operated with an internal volume of 20 µL at a distance of 50 mm from the light source.
@@ 4.2 Deformation of the actuator before and after light-driven operation
The influence of external environmental conditions was investigated by varying the distance between the light source and the actuator (50 mm, 75 mm, and 10 mm), with results presented in Fig. 7D-F (Opteon SF33’s amout is 30 µL). The deformation ratio and plateau bending angle decreased as the distance increased. Notably, the time required to reach the plateau state was longest in the 100 mm group, showing more than a 2.5-fold increase compared to the 50 mm group.
Figure 8: Demonstration of a bio-inspired robot utilizing the proposed actuator in a biomimetic movement driven by light. A: Venus flytrap used as the biomimetic target. B: Conceptual demonstration of the movement, where the robot curves to enclose and retain an object beneath it under light irradiation, and releases it when the light is turning off. C: Actual operation of the robot in the environment, using the same light source as in the experiments. Upon light exposure, the robot captures a miniature ladybug model, protecting and holding it from the light. When irradiation ceases, the robot releases its curvature and frees the enclosed object.
@@ 5 Multi-Finger Robot Using Non-Contact & Light-Driven LIG-Transferred Soft Actuators
To verify the synchronous actuation capability of multiple LIG-transferred soft actuators, we developed a three-fingered robot inspired by the Venus flytrap and evaluated its operation in ambient conditions. The demonstration is shown in Movie S2 and Fig. 8. As illustrated in Fi. 8, upon irradiation from a light source (a lamp in this experiment), the actuators began to bend synchronously, successfully enveloping a ladybug-inspired target object through their coordinated curving motion. When the light source was switched off, each actuator released from its bent state and returned to its original configuration.
@@ 6.1 The operating temperature range control
For the actuator’s internal fluid, we employed Opteon SF33 with a boiling point of 33C. We hypothesize that the operating temperature range can be controlled by varying the boiling point of the encapsulated liquid. For instance, Opteon SF79 and Opteon SF70 have boiling points of 48C and 71C, respectively. Thermal imaging data (Fig. 4B) demonstrates that temperatures approaching 55C can be achieved when LIG is transferred, suggesting that for terrestrial applications, this system could be implemented with liquids having boiling points in the range of 30-55C. However, for operations requiring higher temperature ranges, conventional light sources may prove insufficient. In such cases, the implementation of focusing lenses, high-powered lighting systems, or higher-output lasers is required.
@@ 6.2 Gas leakage
As shown in Fig. 7F, increasing the distance from the light source resulted in a decrease in the bending angle at plateau. This suggests that Opteon SF33 is leaking out due to silicone’s high gas permeability during the slower actuation caused by reduced light intensity. Similar gas leakage phenomena have been discussed in previous studies (?). The current actuator design implements a reusable mechanism where one side is sealed with a clip, allowing multiple refilling of the liquid. However, frequent leakage can limit potential applications. To address this limitation, potential approaches include the incorporation of materials with lower gas permeability and the development of alternative sealing mechanisms.
@@ 6.3 Utilization of ambient environmental light
The light source utilized in this study was a reflector lamp (54W) with a beam angle of 60°, center luminous flux of 630 lm, and beam flux of 160 lm. Assuming uniform light distribution, calculations based on illuminance indicate that the actuator surface, positioned at a distance of 50 mm, receives approximately 60,000 lux (assuming an actuator LIG surface area of 40 $\rm{mm}^{2}$). This illuminance corresponds to approximately 60% of direct summer sunlight (?). While outdoor operation would be dependent on seasonal and geographical factors, the implementation of lens-based light concentration mechanisms could make outdoor applications feasible. Given that typical office environments maintain illuminance levels of approximately 300–1000 lux (?), indoor operation under ambient lighting conditions would be significantly limited. Future investigations will explore localized light energy irradiation using laser sources, aimed at achieving more precise actuator control at a microscale level.
@@ 7 Conclusions
In this paper, we focused on the ocular structure of nocturnal animals that efficiently collect and utilize light in low-irradiation environments. We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization, by incorporating a light-heat conversion element layer called LIG on the light-incident surface. This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, thereby heating the internal low-boiling-point liquid. As a result, the proposed actuator with LIG demonstrated a 54% improvement in response time compared to conventional without-LIG actuators.
While various approaches have successfully enhanced photothermal conversion and thermal conductivity through the incorporation of graphene or metal powders into silicone elastomers, our method explores an alternative approach by installing a LIG layer on the silicone surface. This approach preserves the inherent properties of silicone while utilizing LIG, which can be manufactured under standard environmental conditions, thereby facilitating both the production of photothermal actuators and their diverse applications.
Future research will focus on the versatility of LIG transfer patterns, aiming to develop efficient light-utilizing LIG patterns under natural environmental conditions and design more complex actuator geometries with corresponding LIG patterns. This direction of research aligns with the principles of soft robotics, pursuing the development of soft actuators with enhanced deformation capabilities and environmental adaptability.
@@ Acknowledgement
This work was supported by JSPS KAKENHI Grant Number JP24K21213. I sincerely thank my cats, whose eyes served as a visual reference for the demonstration of the tapetum lucidum structure and whose presence brought inspiration and support throughout the research process.
@@
Maina Sogabe: Conceptualization, Data curation, Investigation, Formal analysis, Methodology,Supervision, Visualization, Writing – original draft, Writing – review & editing
Youhyun Kim: Data curation, Methodology, Formal analysis, Resources, Writing – original draft, Writing – review & editing
Kenji Kawahima: Funding acquisition, Investigation,Formal analysis,Project administration, Supervision, Writing – original draft, Writing – review & editing
@@ 8 Supplementary Material
    Supplementary Figure S1:Changes in temperature icrease at liquid-interface surfaces with directional light irradiation: Thermographic analysis of temperature increase at liquid-contact interfaces comparing two irradiation patterns on LIG-transferred silicone sheets: (A) irradiation from silicone side to LIG side (proposed method) and (B) irradiation from LIG side to silicone side , with (C) showing reference values for silicone-only sheets. The proposed method with irradiation from the silicone side demonstrated higher initial temperature change rate (first 5 seconds after light exposure).
@@
    Supplementary Movie S1 :The actuator’s bending under light irradiation and its bending release upon turning off the light.

### v2

@@ 4.1 Experimental analysis of actuation driven by light On-Off cycling
The experimental setup is illustrated in Fig. 5. During the experiments, the fabricated actuator was positioned 50 mm away from the light source. The light bulb specifications include luminous intensity of 300 cd, total luminous flux of 630 lm, and beam angle of $60^{\circ}$. An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) was used to capture the actuator deformation, and ImageJ (citation) was employed for analyzing the bending angles and deformation measurements.
Figure 6A and C show the bending behavior during light irradiation, the subsequent recovery upon light off, and the corresponding analysis results. When exposed to light, the without-LIG actuator exhibited delayed activation, with a 63% response time of 141.7 seconds and approximately 220 seconds required to reach the maximum bending angle. In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds. It reached maximum bending angle in approximately 90 seconds, showing a 54.1% improvement in responsiveness compared to the without-LIG actuator.
Regarding the recovery phase after light off (Fig 6B and D, the without-LIG actuator showed a 63% response time of 42.7 seconds and required 120 seconds to return to its initial state. The proposed actuator with LIG demonstrated enhanced recovery characteristics with a 63% response time of 22.1 seconds and approximately 60 seconds to return to its initial state, representing a 48.2% improvement in recovery performance compared to the without-LIG actuator.
@@ 4.2 Deformation of the actuator before and after light-driven operation
The influence of external environmental conditions was investigated by varying the distance between the light source and the actuator (50 mm, 75 mm, and 10 mm), with results presented in Fig. 7D-F (Opteon SF33’s amount is 30 µL). The deformation ratio and plateau bending angle decreased as the distance increased. Notably, the time required to reach the plateau state was the longest in the 100 mm group, showing more than a 2.5-fold increase compared to the 50 mm group.
@@ 5 Multi-Finger Robot Using Non-Contact & Light-Driven LIG-Transferred Soft Actuators
To verify the synchronous actuation capability of multiple LIG-transferred soft actuators, we developed a three-fingered robot inspired by the Venus flytrap and evaluated its operation in ambient conditions. The demonstration is shown in Movie S2 and Fig. 8. As illustrated in Fig. 8, upon irradiation from a light source (a lamp in this experiment), the actuators began to bend synchronously, successfully enveloping a ladybug-inspired target object through their coordinated curving motion. When the light source was switched off, each actuator was released from its bent state and returned to its original configuration.
@@ 6.1 The operating temperature range control
We employed Opteon SF33 with a boiling point of 33C for the actuator’s internal fluid. We hypothesize that the operating temperature range can be controlled by varying the boiling point of the encapsulated liquid. For instance, Opteon SF79 and Opteon SF70 have boiling points of 48C and 71C, respectively. Thermal imaging data (Fig. 4B) demonstrates that temperatures approaching 55C can be achieved when LIG is transferred, suggesting that for terrestrial applications, this system could be implemented with liquids having boiling points in the range of 30-55C. However, conventional light sources may prove insufficient for operations requiring higher temperature ranges. In such cases, focusing lenses, high-powered lighting systems, or higher-output lasers are required.
@@ 6.2 Gas leakage
As shown in Fig. 7F, increasing the distance from the light source decreased the bending angle at a plateau. This suggests that Opteon SF33 is leaking out due to silicone’s high gas permeability during the slower actuation caused by reduced light intensity. Similar gas leakage phenomena have been discussed in previous studies[15]. The current actuator design implements a reusable mechanism where one side is sealed with a clip, allowing multiple liquid refilling. However, frequent leakage can limit potential applications. Potential approaches include incorporating materials with lower gas permeability and developing alternative sealing mechanisms to address this limitation.
@@ 6.3 Utilization of ambient environmental light
The light source utilized in this study was a reflector lamp (54W) with a beam angle of $60^{\circ}$, center luminous flux of 630 lm, and beam flux of 160 lm. Assuming uniform light distribution, calculations based on illuminance indicate that the actuator surface, positioned at a distance of 50 mm, receives approximately 60,000 lux (assuming an actuator LIG surface area of 40 $\rm{mm}^{2}$). This illuminance corresponds to approximately 60% of direct summer sunlight[29]. While outdoor operation would depend on seasonal and geographical factors, implementing lens-based light concentration mechanisms could make outdoor applications feasible. Given that typical office environments maintain illuminance levels of approximately 300–1000 lux[30], indoor operation under ambient lighting conditions would be significantly limited. Future investigations will explore localized light energy irradiation using laser sources aimed at achieving more precise actuator control at a microscale level.
@@ 7 Conclusions
This paper focused on the ocular structure of nocturnal animals that efficiently collect and utilize light in low-irradiation environments. We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization. We incorporated a light-heat conversion element layer called LIG on the light-incident surface. This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, heating the internal low-boiling-point liquid. As a result, the proposed actuator with LIG demonstrated a 54% improvement in response time compared to conventional without-LIG actuators.
While various approaches have successfully enhanced photothermal conversion and thermal conductivity by incorporating graphene or metal powders into silicone elastomers, our method explores an alternative approach by installing a LIG layer on the silicone surface. This approach preserves the inherent properties of silicone while utilizing LIG, which can be manufactured under standard environmental conditions, thereby facilitating both the production of photothermal actuators and their diverse applications.
Future research will focus on the versatility of LIG transfer patterns, aiming to develop efficient light-utilizing LIG patterns under natural environmental conditions and design more complex actuator geometries with corresponding LIG patterns. This research direction aligns with the principles of soft robotics, pursuing the development of soft actuators with enhanced deformation capabilities and environmental adaptability.
@@ Acknowledgement
This work was supported by JSPS KAKENHI Grant Number JP24K21213. M.S. sincerely thank my cats, whose eyes served as a visual reference for the demonstration of the tapetum lucidum structure and whose presence inspired and supported me throughout the research process.
@@
Maina Sogabe: Conceptualization, Data curation, Investigation, Formal analysis, Methodology, Supervision, Visualization, Writing-original draft, Writing-review & editing
Youhyun Kim: Data curation, Methodology, Formal analysis, Resources, Writing-original draft, Writing-review & editing
Hiroki Miyazako: Methodology, Writing-review & editing
Kenji Kawashima: Funding acquisition, Investigation, Formal analysis, Project administration, Supervision, Writing-original draft, Writing-review & editing
@@ 8 Supplementary Material
    Supplementary Figure S1
@@
    Supplementary Movie The actuator’s bending under light irradiation and its bending release upon turning off the light.